OpenStudy (pottersheep):

Grade 12 Trig Identities trouble....help please... I'm not sure how to prove that ((sin2x/(1+cos2x)) * (cosx)/(1+cosx) = tan(x/2)

OpenStudy (sirm3d):

try the double angle formulas \[\large \sin 2x=2\sin x \cos x\]and\[\large \cos 2x = 2\cos^2 x -1\]

OpenStudy (pottersheep):

Yep that's what I have tried out...I keep getting stuck though

OpenStudy (pottersheep):

because one side is tan(x/2)

OpenStudy (anonymous):

I've tried it and find what I think is the answer. See if you can find it; I used this page http://www.sosmath.com/trig/douangl/douangl.html

OpenStudy (pottersheep):

Thank you very much I will go take a look